OpenStudy (anonymous):

What is the greatest common factor of 52x^5 and 12x^4?

OpenStudy (anonymous):

52x^5=2^2*13*x^5 12x^4=2^2*3*x^4 GCF=2^2*x^4=4x^4

OpenStudy (anonymous):

These are the answers that I have to choose from 4x^4 4x^5 6x^4

OpenStudy (whpalmer4):

Seen graphically: \(52x^5 = \) 2*2* 13*x*x*x*x*x \(12x^4 = \) 2*2*7* x*x*x*x Collecting the things that appear in both rows: 2*2* x*x*x*x =

OpenStudy (whpalmer4):

@#$#@% it messed up my alignment \[2*2*x*x*x*x=\]

OpenStudy (anonymous):

Thank you this helps!!

OpenStudy (triciaal):

typo the 12x^4 = 2* 2* 3 not 2* 2* 7

OpenStudy (whpalmer4):

oh, yeah, how did I do that? Thanks for pointing it out! Fortunately, that wasn't a common factor :-)

OpenStudy (anonymous):

lolol @whpalmer4

OpenStudy (anonymous):

>.> Stop copying from the internet @LegoMyEgo

OpenStudy (whpalmer4):

I'd rather be good than lucky, but I'll settle for lucky :-)

OpenStudy (anonymous):

Its called using your brain :)

OpenStudy (anonymous):

I saw the exact same answer on yahoo answers which I assume you just copy and pasted

OpenStudy (anonymous):

I am blocking you because I don't like haters.

OpenStudy (anonymous):

When you answered my question, I googled it and could also find the answer on the internet